OpenStudy (anonymous):

Evaluate the following expression. (xy)^0

OpenStudy (abb0t):

1

OpenStudy (shamil98):

Anything to the power of zero is 1, with the exception of zero. You have: \[\huge (xy)^0 = x^0 * y^ 0 =\]

OpenStudy (shamil98):

which is just 1.

OpenStudy (anonymous):

thank you i got some more if you wanna help me?